### v2

# EP 250108a/SN 2025kg: Observations of the most nearby Broad-Line Type Ic Supernova Following an Einstein Probe Fast X-ray Transient

Abstract
With a small sample of fast X-ray transients (FXTs) with multi-wavelength counterparts discovered to date, their progenitors and connections to -ray bursts (GRBs) and supernovae (SNe) remain ambiguous. Here, we present photometric and spectroscopic observations of SN 2025kg, the SN counterpart to the FXT EP 250108a. At , this is the closest known SN discovered following an Einstein Probe (EP) FXT. We show that SN 2025kg’s optical spectra reveal the hallmark features of a broad-lined Type Ic SN. Its light curve evolution and expansion velocities are comparable to those of GRB-SNe, including SN 1998bw, and two past FXT SNe. We present JWST/NIRSpec spectroscopy taken around SN 2025kg’s maximum light, and find weak absorption due to He I m and a broad, unidentified emission feature at 4–4.5 m. Further, we observe broadened H in optical data at 42.5 days that is not detected at other epochs, indicating interaction with H-rich material. From its light curve, we derive a 56Ni mass of 0.2 - 0.6 . Together with our companion paper (Eyles-Ferris et al., 2025), our broadband data are consistent with a trapped or low energy ( ergs) jet-driven explosion from a collapsar with a zero-age main sequence mass of 15–30 . Finally, we show that the sample of EP FXT SNe support past estimates that low-luminosity jets seen through FXTs are more common than successful (GRB) jets, and that similar FXT-like signatures are likely present in at least a few percent of the brightest Ic-BL SNe.
1 Introduction
Though first discovered in the era of sounding rockets (e.g. Cooke, 1976), and highlighted by discoveries with Chandra and XMM-Newton (Jonker et al., 2013; Alp & Larsson, 2020; Quirola-Vásquez et al., 2023), the nature of fast X-ray transients (“FXTs”; non-repeating transients detected in the X-rays) and their connection to the deaths of massive stars and luminous, jetted -ray bursts (GRBs) remains uncertain. By definition, FXTs represent a broad category of transients. In some cases, FXTs could represent the tail of the emission from classical GRBs whose spectra peak at tens to hundreds of keV. FXTs also encompass the population of so-called X-ray Flashes (Heise et al., 2001; Sakamoto et al., 2005), a historical “sub-type” of GRBs defined by a greater ratio of fluence in the X-ray compared to the -ray bands. Most notably, this XRF/FXT group includes the nearby, serendipitous discoveries of Type Ib/c supernovae (SNe), SN 2006aj (Campana et al., 2006; Pian et al., 2006; Soderberg et al., 2006) identified in the long GRB/XRF 060218 and the X-ray-only discovery of the much less luminous SN 2008D (Soderberg et al., 2008). These discoveries showed that at least some FXTs are produced by the core-collapse of massive, stripped-envelope stars. Though sharing similarities with the broad-line Type Ic (Ic-BL) SNe accompanying GRBs (GRB-SNe), a subset of the SNe following these XRFs were times less luminous compared to the prototypical GRB-SN, GRB 980425/SN 1998bw, and one showed absorption due to He (Soderberg et al., 2006; Modjaz et al., 2009), potentially representing a stage in the continuum of massive star explosions (e.g., Galama et al. 1998; Clocchiatti et al. 2011; Soderberg et al. 2008; Alp & Larsson 2020). However, with a limited sample of such events to date, the mapping of GRBs to FXTs, XRFs or core-collapse SNe without high-energy emission continues to be an open question.
Previously, the lack of a dedicated discovery mission meant that the FXT sample was limited to rare, serendipitous detections (e.g., Soderberg et al. 2006, 2008) or archival searches in narrow-field instruments such as Chandra or XMM-Newton (Jonker et al., 2013; Glennie et al., 2015; Bauer et al., 2017; Alp & Larsson, 2020; Novara et al., 2020; Lin et al., 2022; Quirola-Vásquez et al., 2022, 2023). The small (40) population size of FXTs and paucity of multi-wavelength counterparts resulted in large uncertainties on their rates, local environments, and counterpart properties. Uncertainties on these properties inhibit a full understanding of how FXTs fit within the landscape of transients. In particular, many events discovered in the narrow-field searches were of long duration (hundreds to thousands of seconds) compared to GRBs. At present, it is unclear if these long durations represent a selection effect or fundamentally different physical processes (e.g. Levan et al., 2024b). At the same time, the realization of substantial diversity in the progenitors of long GRBs, including events likely produced by compact object mergers (e.g. Rastinejad et al., 2022; Troja et al., 2022; Yang et al., 2022; Levan et al., 2024c), erases the assumption that FXTs arise exclusively via core-collapse.
The 2024 launch of a wide-field soft X-ray monitor onboard the Einstein Probe (EP), opened a new route to discover FXTs and characterize their multi-wavelength counterparts (Yuan et al., 2015, 2022). Critically, EP is able to detect FXTs with its Wide Field X-ray Telescope (WXT; 3600 sq. degree field of view), localize them via WXT to arcminute precision, and subsequently re-point a narrow-field follow-up telescope to obtain arcsecond position. As such, the latency of FXT announcements has reduced from days (and, often, year timescales) to minutes (e.g., Zhang et al. 2024; Li et al. 2025). This speed increase offers the unprecedented ability to locate multi-wavelength counterparts and unveil their origins.
In its first year alone, EP discovered dozens of extragalactic FXTs, revealing a diversity of counterparts and environments. While several FXTs have been associated with GRBs (e.g., Yin et al. 2024; Frederiks et al. 2024; Liu et al. 2025), others have no observed prompt GRB counterpart despite constraints from -ray facilities, prompting speculation that their progenitors are distinct from GRBs produced following core-collapse (e.g., Sun et al. 2024; Bright et al. 2024; van Dalen et al. 2025; Busmann et al. 2025; O’Connor et al. 2025). At least four EP events are known to originate at redshifts (Bochenek et al., 2024; Levan et al., 2024a, b; Liu et al., 2025), representing an exciting new path to exploring the distant Universe but limiting detailed studies of multi-wavelength counterparts. Notably, the counterpart of EP 240414a was localized to a galaxy at and revealed the spectroscopic signatures of a Ic-BL SN. Critically, this event established that at least some EP FXTs originate from the explosions of massive stars (van Dalen et al., 2025; Sun et al., 2024). However, at this redshift, late-time spectroscopic and photometric follow-up of the SN was limited (Sun et al., 2024; van Dalen et al., 2025; Srivastav et al., 2025). More recently, the event EP 250207b was potentially associated with a galaxy at (Levan et al., 2025a). However, at present, no evidence of an SN counterpart has been announced and published observational limits appear to rule out a SN (Yang et al., 2025).
On 8 January 2025 at 12:30:28.34 UT, EP discovered a new FXT, EP 250108a, with a duration of s (Li et al., 2025). Later analysis of Fermi-GBM observations place a conservative, sky-averaged upper limit on associated emission in the 10-1000 keV band of erg cm-2 s-1 at 415 s following the initial trigger (Fermi’s view of the localizaton was occulted by Earth prior to this time; Ravasio et al. 2025). Prompt follow-up observations were not taken, but imaging obtained days after the outburst revealed a blue and rapidly-fading counterpart with a redshift of (Eyles-Ferris et al., 2025). Beginning at days following the initial trigger, the cooling, optical counterpart began to rise in brightness. Spectra taken by the Nordic Optical Telescope (NOT) and Gemini-North at days (where refers to the observer-frame time since the EP trigger) showed the signatures of a Ic-BL SN, dubbed SN 2025kg (see Section 2.2 for further analysis; Levan et al. 2025b; Xu et al. 2025). SN 2025kg offers a unique first opportunity for a detailed comparison between the multi-band light curves and spectroscopic properties of FXT SNe, GRB-SNe and stripped-envelope SNe observed without high-energy counterparts, shedding light on FXT progenitors.
Here, we present optical and infrared observations and analysis of SN 2025kg, including James Webb Space Telescope (JWST) spectroscopy. This work represents a companion paper to Eyles-Ferris et al. (2025) which contains an in-depth analysis of the early ( days post-trigger) observations as well as analysis of the radio and high-energy data. In Section 2 we detail our extensive photometric and spectroscopic campaign to characterize SN 2025kg. In Section 3 we analyze our observations, highlight key features in the JWST spectrum, and compare the properties of SN 2025kg to those of past stripped envelope SNe associated with FXTs, GRBs and discovered without high-energy triggers. In Section 4 we describe our photometric modeling of SN 2025kg. In Section 5, we infer properties of the progenitor from our observations, including its zero-age main sequence (ZAMS) mass, constraints on a He shell due to common envelope mass ejection, and contextualize FXT SNe in the wider population of Ic-BL SNe. Throughout this work, we assume a Planck cosmology, (Planck Collaboration et al., 2020) report all magnitudes in the AB system, and assume an event redshift of (Eyles-Ferris et al., 2025).
2 Observations
2.1 Photometry
Here, we present optical and near-infrared (IR) photometry of SN 2025kg covering days (observed frame; observations at days are presented in Eyles-Ferris et al. 2025). We obtained optical observations with the Alhambra Faint Object Spectrograph and Camera (ALFOSC) on the NOT, the IO:O on the Liverpool Telescope (LT), the Gemini Multi-Object Spectrographs (GMOS) on the Gemini-North and South Telescopes, Pan-STARRS, BlackGEM (for -band description see Groot et al. 2024), the Multi-Object Double-beam Spectrograph (MODS) on the Large Binocular Telescope (LBT), the Goodman spectrograph on the Southern Astrophysical Research (SOAR) Telescope, the Lulin Observatory, MegaCam on the Canada-France-Hawaii Telescope (CFHT), and T80S-Cam on T80S. All optical images were processed using standard techniques, including Gemini DRAGONS (Labrie et al., 2019), a custom python pipeline, POTPyRI111https://github.com/CIERA-Transients/POTPyRI, photpipe (see Rest et al., 2005; Santos et al., 2024, for details), a dedicated LBT data reduction pipeline (Fontana et al., 2014), a modified version of the photometry-sans-frustration package psf (Nicholl et al., 2023), the Elixir pipeline (Magnier & Cuillandre, 2004), a custom built pipeline222https://hdl.handle.net/11296/98q6x4, as well as a combination of standard PyRAF tasks (Science Software Branch at STScI, 2012) and the LACosmic task (van Dokkum, 2001).
We obtain photometry directly on each image using a common set of standard stars drawn from Pan-STARRS (Flewelling et al., 2020). We do not anticipate significant ( mag) contamination from the underlying host galaxy given its faintness relative to the transient as measured in archival Legacy Survey imaging ( mag, mag, mag, mag; Dey et al. 2019). In addition, we obtained near-IR imaging with FLAMINGOS2 on Gemini-South and the MMT and Magellan Infrared Spectrograph (MMIRS) on the MMT. We process all near-IR images using DRAGONS and POTPyRI and obtain photometry using a common set of standard stars from the Two Micron All Sky Survey (2MASS; Skrutskie et al. 2006).
We provide details for all photometric programs in Appendix Table 1. We further incorporate detections from the GCNs (Ror et al., 2025; Song et al., 2025; Zou et al., 2025) and ATLAS forced photometry (-bands; Tonry et al. 2018). We present all photometry in AB magnitudes in Appendix Table 2 and show our observations in Figure 1.
2.2 Spectroscopy
We present thirteen optical spectra of SN 2025kg taken over days. Spectroscopy prior to the emergence of SN features ( days) is presented and analyzed in our companion paper (Eyles-Ferris et al., 2025). Our optical spectroscopy was obtained with GMOS on the Gemini-North and South Telescopes (Program IDs GN-2024B-Q-107, GN-2024B-Q-105, GS-2024B-Q-105, GS-2025A-Q-107, P.I.s Rastinejad, Chambers), OSIRIS+ on the Gran Telescopio Canarias (Program ID GTC1-24ITP; P.I. Jonker), the Multi Unit Spectroscopic Explorer (MUSE) on the Very Large Telescope (VLT; Program ID 111.259Q.001, P.I. Jonker), and the Low Resolution Imaging Spectrometer (LRIS) on Keck Observatory (P.I.s Harrison, Prochaska). Details of each spectroscopic set up are logged in Appendix Table 3.
We reduce Keck and Gemini spectroscopy using PypeIt (Prochaska et al., 2020). To account for instrumental flexure, sub-pixel shifts in the position of the [O I] 6300 Å sky emission line were measured and applied as corrections. We reduce GTC spectra using Molly to correct for the Earth’s motion relative to the target and observations of spectrophotometric standard stars taken the same night for flux calibration. We reduce the VLT spectrum using ESO-reflex (Freudling et al., 2013). We process the LBT spectrum using the Spectroscopic Interactive Pipeline and Graphical Interface (SIPGI) tool (Gargiulo et al., 2022) and perform wavelength calibration using arc lamp frames. To obtain flux-calibrated LBT spectrum, we applied the sensitivity function derived from the spectro-photometric standard star Feige 34. We correct all spectra for Galactic extinction in the direction of the burst ( mag; Schlafly & Finkbeiner 2011) and show the optical spectroscopic series in Figure 2.
In each spectrum, we identify the distinct signatures of Type Ic-BL SNe. For instance, we observe at least two broad absorption features around the expected locations of Fe II (or a blend of the Fe triplet) and Si (Figure 2) and no prominent absorption around the optical features of H or He. In several of the spectra we observe narrow emission lines, including those at the rest-frame locations of H, H, and [OIII], which we attribute to the underlying host galaxy. To compare SN 2025kg to other SNe, we utilize the spectral template matching tool, gelato (Harutyunyan et al., 2008), which matches an input spectrum to a large bank of observed SNe spectra at a range of phases. For the SN 2025kg spectrum observed at days gelato finds a strong match to the Type Ic-BL SN 2006aj (associated with GRB/XRF 060218) at 0.8 days past peak provides the strongest fit. For the SN 2025kg spectrum observed at days, gelato finds strong matches to several Type Ic-BL SNe, including SNe 1997ef, 1998bw, 2002ap and 2006aj. This analysis definitively confirms our classification of SN 2025kg as a Type Ic-BL SN.
In the GMOS spectrum at days (Figure 3) we observe a broadened H emission feature, distinct from the narrow H emission observed in previous epochs (Figure 2). We inspected the individual four 2D spectral frames of the days spectrum and observe evidence for excess emission associated with this broadened feature in each. We do not see significant broadened H emission in the previous ( days) or subsequent ( day) spectra. We discuss this feature in greater detail in Section 3.2.
In addition, we obtained two near-IR spectra of SN 2025kg. At days post-burst we obtained a cross-dispersed spectrum covering m with the Gemini Near-InfraRed Spectrograph (GNIRS) on Gemini-North (Program ID GN-2024B-Q-107, P.I. Rastinejad). These data were processed in Pypeit with additional post-processing performed following methods described in Tinyanont et al. (2024). Due to low signal-to-noise over the majority of the bandpass, we consider only the portion m usable. Further, at days, we obtained fixed slit spectroscopy with NIRSpec on JWST (Program 6133; P.I. Gompertz) using the prism (covering 0.5-5 microns) at low resolution, and with an exposure time of 6302 seconds. At the time of the JWST spectroscopy the source is point-like and several magnitudes brighter than the anticipated magnitude of its compact underlying host galaxy. Therefore, we utilize the default pipeline reduction and 1D extraction of the NIRSpec observations.
3 Analysis & Comparison to Past Events
3.1 SN Light Curves
To place SN 2025kg in context of past GRB-SNe, FXT-SNe, and other stripped-envelope SNe, we compare its light curve to those of relevant past events. We include eleven events with X-ray or GRB counterparts: SN 1998bw (GRB 980425; Clocchiatti et al. 2011), SN 2003dh (GRB 030329; Matheson et al. 2003), SN 2006aj (GRB/XRF 060218; Mirabal et al. 2006; Soderberg et al. 2006), SN 2008D (XRF 080109; Soderberg et al. 2008) SN 2010bh (GRB/XRF 100316D; Cano et al. 2011; Olivares E. et al. 2012), SN 2011kl (GRB 111209A; Levan et al. 2014; Greiner et al. 2015) SN 2013cq (GRB 130427A; Perley et al. 2014), SN 2016jca (GRB 161219B; Cano et al. 2017), SN 2017iuk (GRB 171205A; Izzo et al. 2019), SN 2019oyw (GRB 190829A; Hu et al. 2021; Rastinejad et al. 2024), and SN 2024gsa (EP 240414a; Srivastav et al. 2025). For events without high-energy counterparts, we show a subset of Type Ib/c and Ic-BL SNe from the literature (Foley et al., 2003; Drout et al., 2011; Izzo et al., 2020; Ho et al., 2020), which we normalize in time to the peak of SN 2025kg in each respective filter. We show all light curves in the nearest rest-frame filter using their known redshifts (Figure 4).
This comparison (Figure 4) highlights that SN 2025kg’s luminosity is comparable to most GRB-SNe, particularly SN 1998bw, and distinct from SNe observed without high-energy counterparts. Amongst the SNe associated with FXTs, the closest analogs of SN 2025kg are SN 2006aj and SN 2024gsa, the latter of which is also associated with an EP FXT. The optical light curve of SN 2024gsa is marked with a large flare between days (e.g., Srivastav et al. 2025; van Dalen et al. 2025), which we do not observe in early observations of SN 2025kg (Eyles-Ferris et al., 2025).
3.2 SN 2025kg Spectra & Broadened H at days
We next attempt to compare our spectra of SN 2025kg to past GRB-SNe, FXT SNe, Ic-BL SNe without a high-energy trigger, and AT 2018cow, the latter of which is a benchmark event showing broadened H emission (Prentice et al., 2018; Margutti et al., 2019; Perley et al., 2019). We download spectra from WISeREP (Yaron & Gal-Yam, 2012) of the prototypical GRB-SN, SN 1998bw (Patat et al., 2001), three SNe associated with FXTs, SN 2006aj (Pian et al., 2006), SN 2008D (Modjaz et al., 2006, 2014) and SN 2010bh (Bufano et al., 2012) and SN 2020bvc (Hiramatsu et al., 2020). Further, we incorporate spectra of the luminous fast blue optical transient, AT 2018cow, in which broadened H was prominently observed starting at 15 days following the event’s detection (e.g., Margutti et al. 2019; Perley et al. 2019; Xiang et al. 2021).
In Figure 5 we show the spectra for GRB-SNe and FXT SNe within a week of the SN peak ( days) and the spectrum of SN 2020bvc taken 1.9 days after the explosion (no later spectra are available on WISEReP). Spectra of the Type Ib SN 2008D show prominent He I 4471, 5876, 6678, 7061 Å absorption lines (Modjaz et al., 2009), which we do not observe in SN 2025kg. Considering the remaining spectra, at both epochs, SN 2010bh provides a poorer fit to SN 2025kg than the SN 2006aj and SN 1998bw, which we attribute in part to its larger ejecta velocities (see Section 3.3; Chornock et al. 2010). Similar to their light curve properties (Section 3.1) SN 1998bw and SN 2006aj are spectroscopically comparable objects to SN 2025kg (Section 2.2). SN 2020bvc also shows a comparable shape to SN 2025kg, in keeping with comparisons between the two objects in our companion paper (Eyles-Ferris et al., 2025)
Turning to the broadened H emission, we compare our SN 2025kg spectrum at days to those of SN 1998bw, SN 2010bh and AT 2018cow at days in the region of H (Figure 6). Spectra of SN 2006aj beyond 20 days of the initial high-energy trigger are not available on WISeREP. Compared to the spectrum of SN 2025kg binned at a resolution of 40 Å, the broadened emission feature is stronger and more narrow in AT 2018cow and absent or very weak in SN 2010bh. Though a comparable bump in the spectrum of SN 1998bw is observed, we note that this likely due to two absorption features on either side of H. Further, in SN 1998bw the spectral shape in this region is is broadly consistent across several weeks of observations. In contrast, in SN 2025kg the continuum is fairly flat over Åat most epochs, but exhibits broad H at one single epoch (Figure 3). Thus, we conclude that absorption features cannot explain the broadened emission feature in SN 2025kg, and that its most likely interpretation is that it arises from H interaction.
Whether this feature is common in FXT SNe but not GRB-SNe, or if SN 2025kg is an outlier amongst other FXT SNe remains to be seen. Indeed, since the H appears only in one spectrum it may be a transient feature that could occur more frequently in other Ic-BL SNe but is often missed because it appears only at later times, and is only visible for a short period. One potential explanation for this feature in SN 2025kg is interaction with an extended H shell of circumstellar material (CSM), as was considered for AT 2018cow (Margutti et al., 2019; Perley et al., 2019). However, in AT 2018cow the broadened H signature was stronger compared to the continuum and persisted over several weeks, and was generally narrower, whereas in SN 2025kg the feature is weaker and absent in the spectrum at days (Figure 2). Alternatively, this could be a signature of a binary companion, as is observed in a small (1-5%) fraction of Type Ia SNe (Maguire et al., 2016; Tucker et al., 2020). A larger sample of spectra at days is critical to exploring the presence and temporal evolution of broadened H in future FXT SNe. We further discuss implications of this feature in Section 5.3.
3.3 Broad Line Velocity Measurements
We compute the ejecta expansion velocity using the Doppler shift of the absorption minima ascribed to Fe II Åand Si Å, which are well-characterized in GRB-SNe (e.g., Chornock et al. 2010; Bufano et al. 2012). For each spectrum, we fit the wavelength region around the absorption feature with a linear continuum and Gaussian profile using scipy.curvefit, and determine the velocity using the absorption minima’s shift relative to the rest-frame wavelength. We show the results from our fitting over time and compare to values for several GRB-SNe and an SN Ic-BL (Chornock et al., 2010) in Figure 7.
We derive the velocity evolution for both Fe II and Si . Though slight discrepancies between the velocities are present at several epochs, we note that measurements of expansion velocity using different lines are known to produce inconsistent values, in part due the locations of each elements within different layers of the ejecta (Modjaz et al., 2016; Finneran & Martin-Carrillo, 2024). We next compare to the Si velocities of SNe 1998bw, 2002ap, 2006aj and 2010bh (Chornock et al., 2010) and the only other confirmed EP SN, SN 2024gsa (EP 240124A; Sun et al. 2024; van Dalen et al. 2025), for which one measurement at days is available (Figure 7, bottom panel; Sun et al. 2024). At this epoch, the expansion velocities of SN 2025kg and SN 2024gsa are comparable. Compared to the wider sample of events, SN 2025kg appears most closely related to SN 2006aj, although the temporal coverage is not equivalent. SN 2025kg has larger measured velocities compared to the Ic-BL SN 2002ap and, generally, lower velocities compared to SN 1998bw and SN 2010bh.
3.4 Near-IR Evidence for Helium
Motivated by questions of the unknown stellar progenitors of FXTs, we next investigate to what degree He is present in the spectra of SN 2025kg. In our optical spectra, we do not observe prominent He I 4471, 5876, 6678, 7061 Å absorption lines, as were observed in SN 2008D (Figure 5; Soderberg et al. 2008). However, as He produces stronger signatures in the near-IR compared to the optical, we focus our search on the He I lines m in the near-IR spectra. He I is a prominent feature that dominates the 1 m region in Type Ib SNe, and may be blended with C I m and Mg II m lines in Type Ic/Ic-BL SNe (e.g., Shahbandeh et al. 2022; Tinyanont et al. 2024). Though it is well-established that the He I m line is a more dominant line compared to He I m, the redder He I line has a greater offset from other potential absorption features and, thus, may be easier to distinguish as He I (Shahbandeh et al., 2022; Tinyanont et al., 2024). In the GNIRS spectrum, we do not observe any significant absorption features in the 1 m regime, though we note that the majority of the spectrum has a low signal-to-noise.
Turning to the JWST spectrum, we observe a prominent, broadened absorption feature 1 m and a shallower absorption feature 2 m (Figure 8). We first explore if the 2 m feature can be explained with only C I m and Mg II m or a combination of the two (i.e., no He I), using a Bayesian toolkit to model blended absorption features from multiple species with Gaussian components (Liu et al., 2023)333https://github.com/slowdivePTG/BayeSpecFit. find velocities in the range of 24,000 - 41,000 km s-1, 10,000 km s-1 faster than the Fe II and Si expansion velocities at a similar epoch (Figure 7). This leads us to disfavor that this feature can be explained by C I and Mg II alone. A two-Gaussian fit of He I and Mg II provide somewhat high (relative to those of Fe II and Si) velocities of 20,000 - 22,000 km s-1, and favors a larger contribution from Mg II than He I. Overall, we conclude that the 2 m feature is likely to be at least partially due to He I absorption, and do not rule out marginal blending with Mg II.
We next consider whether the 1 m feature is due to He I absorption. If both this feature and the 2 m feature are dominated by He I, the expansion and dispersion velocities of both features would be consistent and the equivalent width ratio, EW(He I m)/EW(He I m) (Lucy, 1991). To test these, we first assume this feature is due to He I only, and derive an expansion velocity of 19,000 km s-1, consistent with modeling of the light curve (Section 4) a dispersion velocity of 9,000 km s-1, and a pseudo-EW ratio of 1.47, within the expected range. The larger dispersion velocity of the the 1 m feature relative to the 2 m feature (Figure 9) suggests that the 1 m feature is likely a blend of He I, C I and/or Mg II.
Finally, we simultaneously fit the 1 and 2 m features assuming both could be a combination of the He I lines m and the C I lines m (Figure 9). This fit finds He dominates the 2 m feature and is blended with C I in the 1 m feature. This fit describes the data well and gives a pseudo-EW ratio between the two best-fit He I absorption lines. We conclude that there is likely a small amount of He in the ejecta of SN 2025kg, though a precise estimate is outside the scope of the current work. Blending of He I and another feature is supported by a higher-resolution near-IR spectrum of SN 1998bw spectrum showing two “sub-features” within the 1 m feature (Figure 9). Based on the absence of He I lines in the optical (Figures 2 and 5) we place a conservative upper limit on He in the ejecta of (see, e.g., models for SN Ib/c optical spectra in Dessart et al., 2020).
In Figure 9 we show the SN 2025kg JWST spectrum and a near-IR SN 1998bw spectrum (8 days post-peak; Patat et al. 2001) in velocity space relative to the He I and Mg II features. Both spectra show evidence for absorption around the 1 and 2 m features, with the 1 m feature being most prominent. Near-IR spectra of SN 1998bw were taken at 8, 33 and 51 days post-peak, and each shows a prominent feature near 1 m ascribed to He I (Patat et al., 2001) which was not observed in a near-IR spectrum of SN 2010bh (Chornock et al., 2010). Further, in the SN 1998bw spectrum at +51 days post-peak Patat et al. (2001) also claim the detection of He I m and state this line is weakly detected in the spectra at +8 and +33 days post-peak (Patat et al., 2001). We observe a similar shape 2 m in the spectra of SN 2025kg and SN 1998bw around peak (Figure 9). Altogether, this comparison demonstrates that small amounts of He are likely common in GRB-SNe and FXT SNe, but are not typically observed due to a paucity of near-IR spectroscopic coverage.
3.5 4-4.5 m Feature
We next turn to the broad emission feature centered at a rest-frame wavelength of m in JWST/NIRSpec spectrum taken around SN 2025kg’s peak (Figure 8). Here, we discuss several potential explanations for this feature, but ultimately determine that a larger sample of observed events is critical to establishing its origin.
First, we investigate the possibility that this feature is a signature of -process nucleosynthesis. Infrared excesses have been suggested as possible evidence of -process, which may occur in the accretion disk after the collapse of massive, rapidly-rotating stars, that are thought to be the progenitors of GRBs and some FXTs (e.g., Siegel et al. 2019; Zenati et al. 2020). To test this possibility, we explore if the m feature is consistent with a thermal signature. We find that this condition is not satisfied, as the feature is too narrow to be well fit with a Planck function. Further, the peak wavelength implies a temperature of K, from which we infer a radius of cm to match the luminosity. This would require an average expansion velocity of c over the first two rest-frame weeks, which is incompatible with both observations (Section 3.3) and theoretical expectations.
Second, we consider the possibility that this feature is due to dust. To investigate this, we consider comparable spectroscopic observations of SNe at mid-IR wavelengths to investigate their timescales for dust emergence, which are thus far limited in the literature. JWST NIRSpec and MIRI spectroscopic observations of the Type IIP SN 2022acko show no evidence for dust at 50 days following the explosion (Shahbandeh et al., 2024). In SN 1987A, near-IR spectra cover the range m – m (Meikle et al., 1989). The 4 m region is dominated by a narrow emission feature that is attributed to Brackett ( m). H features are not expected in a Ic-BL SN like SN 2025kg, although we note a narrow excess at this wavelength in our spectrum (Figure 8), which may be emission from the underlying host galaxy. In SN 1987A, at around 260 days post-peak a comparable broad underlying feature emerged and was tentatively suggested to be due to an overtone of silicon monoxide (SiO) (Meikle et al., 1989). Both of these epochs are significantly later than our JWST spectrum of SN 2025kg, which was taken at rest-frame days after the FXT trigger. Additionally, at this point, the ejecta temperature is likely too high for dust molecules to form, supporting that this feature is not SiO. Further, spectral synthesis modeling indicates that the SiO overtone emission emerges at days following core-collapse in Type Ic-BL SNe (Liljegren et al., 2023).
Finally, the feature may be the result of nebular Fe lines. This interpretation would require regions of the ejecta to become optically thin, which may be possible with high expansion velocities. Expansion velocities of c were inferred at early time in Eyles-Ferris et al. (2025), which may be sufficient to produce this signature. Looking forward, the assembly of a sample of mid-IR spectra of stripped envelope SNe or other FXTs will be important to understand how ubiquitous this feature is, and shed light on its physical origin.
3.6 Comparison of SN and High-Energy Properties
In our companion paper, we show that EP 250108a/SN 2025kg’s prompt FXT, early ( days) optical light curve, and radio upper limits are consistent with a trapped or weak jet from a collapsar (Eyles-Ferris et al., 2025). Motivated to explore if this event is an typical FXT SN, we consider any potential correlations between the high-energy and SN properties of FXTs (Figure 10).
In addition to EP 250108a/SN 2025kg, we include the events SNe 1998bw, 2006aj, 2008D, 2010bh, and 2024gsa in this comparison. We utilize the prompt X-ray (2-26 keV) emission of SN 1998bw detected by BeppoSAX (Galama et al., 1998; Pian et al., 2000), X-ray detections by Swift-XRT of SN 2006aj, SN 2008D, and SN 2010bh pulled from UKSSDC (Evans et al., 2007, 2009), and the prompt X-ray emission of SN 2024gsa and SN 2025kg observed by EP (Li et al., 2025; Sun et al., 2024). We caution that these properties are drawn from a variety of instruments, each with different energy ranges and slewing strategies. However, we note that these comparisons span several orders of magnitude, mitigating some biases. We mark the different instruments in Figure 10. For SN properties, we incorporate SN peak magnitudes derived from the light curves presented in Sections 3.1 and Si velocities presented in Section 3.3.
Figure 10 highlights that SN 2025kg, SN 1998bw and SN 2006aj occupy a similar region in the parameter space of peak SN and peak X-ray luminosity (see also Figure 13 of Eyles-Ferris et al. (2025)). We observe a potential trend in these properties for the objects outside of this cluster, though a significantly larger sample is needed to test this trend further. SN 2025kg and SN 2006aj are also comparable in their observed FXT duration. However, we caution that the duration of EP events, in particular, is often set by instrumental rather than astrophysical effects and the duration of EP 250108a is a preliminary value (Li et al., 2025). Of the Swift events, there is a potential trend between SN peak brightness and observed FXT duration. Notably, EP 250108a/SN 2025kg has the longest X-ray duration of the events considered here.
4 Radioactive Decay SN Modeling
With this extensive data set, we are motivated to constrain the physical properties of SN 2025kg, which enable comparisons to previous SN and possible properties of the progenitor. Thus, we fit the light curve for SN 2025kg using Redback (Sarin et al., 2024). We fit the light curve using two models: first, a one-zone “Arnett” model (Arnett, 1982) and second, a model that accounts for mixing of 56Ni into the outer ejecta.
For both models, we fit the entire light curve at days assuming a Gaussian likelihood. For each photometric point, we include a systematic error of 0.15 mag added in quadrature to the statistical errors to capture any discrepancies caused by differences in photometric reduction, filter transmission curves, or in the model itself. We explore the model parameter space using Bayesian Inference via the Pymultinest nested sampler (Buchner et al., 2014) through the Bilby library (Ashton et al., 2019).
4.1 One-zone Model
We begin with a one-zone “Arnett” model that assumes a one-zone, homologously-expanding ejecta powered by a centrally-located heating term (i.e., the 56Ni) with a gray ejecta opacity that is constant in time. This model is commonly invoked to describe the physical properties of SNe. We consider model parameters for the one-zone model here: ejecta mass, ; fraction of 56Ni in the ejecta, ; ejecta velocity, ; gray opacity, ; -ray opacity, ; temperature at which the photosphere begins to recede, ; and host extinction in magnitudes, . For the final parameter, we assume a standard Fitzpatrick extinction law with an (Fitzpatrick, 1999). We use broad uninformative priors on each parameter.
Our model provides a strong fit to the data. We infer of ejecta with a 56Ni mass of and an initial ejecta velocity of km s-1. Our estimate of ejecta mass and energy is consistent with median values inferred from Type Ic-BL SNe (Taddia et al., 2019; Srinivasaragavan et al., 2024) using a similar model. However, our estimate of the 56Ni is marginally inconsistent with the Type Ic-BL SNe medial values of Taddia et al. (2019) (at ), which could be due to differences in assumptions about -ray opacities or hinting towards a limitation with the one-zone model.
4.2 Nickel Mixing Model
There are clear limitations with one-zone model in its assumption of a centrally-located 56Ni and constant opacity. In reality, the opacity is almost certainly not gray or constant with time (e.g., Niblett et al., 2025) and 56Ni is likely mixed out to large radii due to Rayleigh-Taylor or jet-driven instabilities (e.g., Khatami & Kasen, 2019; Reichert et al., 2023). The latter effect significantly impacts the overall bolometric light curve. In particular, the classical “Arnett-like” one-zone model often over-predicting the quantity of 56Ni compared to models with mixing (e.g., Bersten et al., 2014) for two reasons. First, the central heating term is forced to overcome adiabatic losses compared to heating from ejecta at larger radii. Second, energy through heating has to diffuse through more ejecta. Therefore, our 56Ni mass measurement above is likely an overestimate.
To better ascertain the quantity 56Ni and to also explore the impact of mixing on our light curve fits and estimated parameters, we also fit a 56Ni-mixing model. We outline the details of this model below, but detailed derivations and comparison to the one-zone model will be presented in a forthcoming publication (Sarin et al., in prep). This derivation is largely inspired by the kilonova model outlined in Metzger (2019). We begin by assuming a power-law distribution of masses under the assumption of homologous expansion
where is an arbitrary coefficient such that the sum of all layers is the total ejecta mass, , is the minimum velocity of the outflow, and describes the mass in each ‘layer’. The number of layers is arbitrary, provided that they are sufficiently thin to accurately capture the dynamical evolution of the ejecta. A better description of SN ejecta is likely a broken power law (e.g., Matzner & McKee, 1999). However, the deviations in light curve for different density profiles are smaller than the impact of mixing and other uncertain physics such as -ray leakage (e.g., Sollerman et al., 1998). The thermal energy of each layer evolves following the first law of thermodynamics,
where describes the energy input into each layer from radioactive decay, the second term on the right-hand-side describes losses due to PdV expansion and describes radiative losses from each layer, which needs to account for energy losses due to the light crossing times (e.g., Metzger, 2019). These equations can be numerically solved while distributing the total 56Ni into layers (following the same distribution as the original mass distribution) up to a certain mass layer set by an additional parameter to capture nickel mixing. The photospheric radius is set as the radius of the mass shell at which the , which starts to recede back into the ejecta when the temperature drops below . We further improve the physics of the model by incorporating a temperature-dependent opacity following,
where is the effective opacity, which smoothly transitions from the maximum opacity, to the minimum opacity, as the temperature starts to approach . This functional form and are chosen to replicate the opacity evolution in detailed numerical simulations (e.g., Nagy, 2018).
We fit this model to the observed light curve as we did in Section 4.1. Our fits with this model are shown in Figure 1 highlighting broad agreement with the data. Compared to the one-zone model, our estimated parameters better match intuition and expectations. In particular, we infer a total ejecta mass of with a 56Ni mass of . We also find evidence for significant mixing, with , that is, 56Ni is distributed out into of the mass layers. The lower 56Ni estimate compared to our one-zone model is driven by the high mixing, which reduces the need for a larger central 56Ni as the mixed material has to overcome less adiabatic losses and diffuse through less ejecta.
Though this model offers several advantages compared to the one described in Section 4, there are still uncertainties, such as the assumed mass profile, therefore, we conservatively use the 56Ni estimate from both models. These masses and mixing provide some strong hints into the progenitor, as we discuss in Section 5. Given the limits of both models, we adopt a conservative range based on both for the 56Ni mass of
5 Discussion
Our analysis of the light curves and spectra of SN 2025kg place strong constraints on the nature of the progenitor. In our companion paper (Eyles-Ferris et al., 2025) we argue that the engine behind this transient is jet-driven, supported by shock break-out model fits to the prompt X-ray emission, upper limits on radio emission and the blue and rapidly-fading light curve prior to the onset of the SN. In this section, we combine the inferences of both papers to discuss constraints on the progenitor and implications for the population of FXTs from collapsars.
5.1 Constraints on the ZAMS Mass of the Progenitor
Here, we determine the ZAMS mass of the progenitor by combining the inferred masses on the ejecta yield, the 56Ni yield (which dictates the disk mass) and the compact remnant mass. To do this, we set the mass of the star at explosion to the C/O core mass as SN 2025kg shows evidence for stripping of its H and most of its He layers; Section 2.2. We assume that the C/O core mass will be equivalent to the summation of the ejecta mass, the disk mass and the compact object remnant mass. For the first component, we use the results from our light curve fits, which indicate a total ejecta mass of with a 56Ni mass of in SN 2025kg (Section 4). We explore two scenarios to determine the latter two components: a black hole accretion disk jet model and a model invoking a neutron star accretion disk. If the remnant is a black hole, we assume the accretion disk forms around a black hole. If it is a neutron star, we assume an initial remnant mass of .
The large amount of 56Ni ejecta mass required to explain the SN light curve peak (Section 4) indicates that a considerable mass must be processed through an accretion disk. In a jet-driven explosion, the 56Ni is primarily produced in the high-temperature disk and ejected through a disk wind (Surman et al., 2006). Typically models predict that % of the disk is lost through a wind (e.g Aktar et al., 2017; Siegel & Metzger, 2017). This means that the mass that flows into the disk must be greater than times that of the 56Ni ejecta mass. To eject of 56Ni, roughly would have to be processed through a disk. Combined with the non-nickel ejecta mass and the compact remnant mass, this corresponds to a C/O core of for a neutron star accretion disk model or for a black hole accretion disk model. Using the C/O star masses from low-metallicity stars simulated by (Woosley et al., 2002), the C/O core masses correspond to a ZAMS progenitor masses of if the compact remnant is a neutron star and if the compact remnant is a black hole. These mass ranges are consistent with our current expectations for the progenitors of neutron stars and black holes respectively (Fryer, 1999). The uncertainty in the 56Ni mass dominates the wide range of our predicted progenitor masses. A late-time observation of the light curve will help constrain the 56Ni mass and place stronger constraints on the progenitor’s ZAMS mass. Finally, we note that our above estimates are based on single star models with a fixed set of parameters for stellar evolution (e.g. mixing; Woosley et al. 2002). Models that include the effects of binary mass transfer and ejection are likely to alter or broaden the predicted range of progenitors. However, these results provide a first pass of the stellar conditions.
5.2 Constraints on a He Shell
We next consider constraints on a He shell around the progenitor of EP 250108a/SN 2025kg from our observations. The early optical light curve of the counterpart of EP 250108a ( days; Figure 1 and Eyles-Ferris et al. (2025)) is marginally consistent with interaction of a shell at a distance of . This distance is consistent with being produced by a common envelope mass ejection of a He shell, and just below the lower radii predicted for He shell mass ejection from a common envelope if it occurs at C ignition (Fryer et al., 2024). This distance is also consistent with a He merger event (Fryer & Woosley, 1998; Grichener, 2025).
Using the He shell models of Woosley et al. (2002) models, we predict a large amount of He in this shell (). Notably, this is inconsistent with our upper limit on the He mass of from the optical spectra (Section 3.4). This He mass could be reduced if significant Wolf-Rayet mass loss occurs prior to the final mass ejection episode that produces the shell at . Alternatively, enhanced mixing (e.g. Frey et al., 2013) could reduce the size of the He shell. In conclusion, for our models to fit the data presented in Eyles-Ferris et al. (2025) and this work, either considerable wind mass-loss or stellar mixing is required.
5.3 Implications of Hydrogen Observed at days
In Section 2.2 we show that broadened H is observed at days, or when the shock reaches . Similarly, the Type I superluminous SN iPTF13ehe showed a similar late-time H feature initially observed in spectra at +251 days from peak light, which was interpreted as coming from a detached shell of H around 1016 cm from the progenitor star (Yan et al., 2017).
Here, we briefly consider several potential explanations for this feature. We find it unlikely that the H is due to the progenitor star’s wind, as models robustly predict that a Wolf-Rayet wind would evacuate the H in a region with radial extent of 0.1-1 pc (Weaver et al., 1977). One potential explanation could be that the H is from a common envelope mass ejection phase with a sufficient asymmetry to produce clumps of H-rich material that could persist despite the strong Wolf-Rayet wind. Alternatively, the H signature could be produced through interaction with a companion star. However, such a wide separation would argue that this companion star had no impact on the stellar evolution of the collapsing star. To ensure that the supernova blastwave sweeps up enough H to produce the H line, this companion would have to be evolved, either in a giant or supergiant phase (Hirai et al., 2018, 2020; Hirai, 2023). Since we are likely to require a close binary to explain the rotation and loss of the He shell, this would then imply that the progenitor of SN 2025kg is in a triple system.
5.4 A Growing Population of FXT-SNe & Connections to GRB-SNe and Ic-BL SNe
Early works on XRF/FXTs with SNe found that the volumetric rates of such events would be at least an order of magnitude higher than those of collapsar GRBs (Soderberg et al., 2008). EP 250108a/SN 2025kg joins a population of three FXT-SNe (EP 240414a/SN 2024gsa, EP 250304a; van Dalen et al. 2025; Sun et al. 2024; Izzo et al. 2025) discovered in just the first year of EP operations. Of those with published SNe light curves and spectra (EP 240414a/SN 2024gsa and now EP 250108a/SN 2025kg) these events are comparable in their peak optical luminosities, optical spectra, and Si 6355 Å absorption velocities (Sections 3.1-3.3) to GRB-SNe and the XRF/FXT 060218/SN 2006aj. Further, in Section 3.6, we demonstrated that the peak X-ray luminosities of GRB 980425, XRF/FXT 060218, and EP 250108a are also comparable. Taken together, these support a causal link between the progenitors and mechanisms driving GRBs and FXTs with SNe.
The three recent EP FXT detections and the rates of GRB-SNe discovered over the last 20 years (1 per year; e.g., Dainotti et al. 2022) support earlier calculations (Soderberg et al., 2008) that FXT are significantly more common than GRB-SNe. While a robust calculation of the relative rates of FXT SNe and GRB-SNe is outside the scope of this work, we note that the three FXT SNe discovered in the first year of EP operations were at and and detected by an instrument covering of the sky. From this, we derive an estimate of the intrinsic volumetric rates of at least several tens Gpc-3 yr-1, 1-2 orders of magnitude larger than the rate of on-axis luminous GRBs (Sun et al., 2015).
In our companion paper, Eyles-Ferris et al. (2025) we demonstrate that EP 250108a/SN 2025kg is consistent with being driven by a collapsar-powered jet that either fails to break out of the dense circumstellar material or has an energy weaker than ergs. Taken together with the rates, this indicates that trapped or weak jets are more common than the successful jets that produce GRBs. The lack of radio and X-ray detections at later times following EP 250108a/SN 2025kg (Eyles-Ferris et al., 2025) implies that observing any relativistic material may be challenging from these FXT-producing jets, in keeping with late-time radio surveys of Ic-BL SNe observed without high-energy counterparts (e.g., Corsi et al. 2016). Zooming out, it is like that FXT-producing, “trapped” jets accompany a substantially larger fraction of Ic-BL SNe than those with successful jets and GRBs, representing at least a few percent of the Ic-BL SNe population.
Finally, we note that the SNe of two past FXT events, SN 2008D and SN 2010bh, are 5-100 times less luminous than FXT SNe 2006aj, 2024gsa, and 2025kg, and GRB-SNe and, in the former’s case, show prominent He absorption (Figure 5; Soderberg et al. 2008; Modjaz et al. 2009). The lower optical luminosities of these SNe are echoed in the lower peak X-ray luminosities of their FXTs (Figure 10). Future EP detections of FXT SNe will reveal both the rates of such events and the full span in stellar progenitors of FXT SNe.
6 Conclusion
Along with our companion paper (Eyles-Ferris et al., 2025), we have presented the most detailed dataset to date of an SN accompanying an EP FXT. Our main conclusions are as follows:
Optical spectra of SN 2025kg are characterized by broad absorption features due to Fe II and Si and do not show obvious emission due to He I, leading to a Type Ic-BL classification.
SN 2025kg is a close analog of GRB-SNe, particularly SN 1998bw, and the FXT SNe, SN 2006aj and SN 2024gsa, in terms of its peak luminosity, expansion velocities, and spectral evolution. We also find a strong match between SN 2025kg and the Ic-BL SN 2020bvc, in keeping with comparisons to the early light curve (Eyles-Ferris et al., 2025).
The 1 and 2 m features in the JWST spectrum taken around maximum light are at least partially due to He I absorption. From optical spectra, we conservatively conclude that the mass of He in the ejecta is .
We observe a broad feature at 4-4.5 m in the JWST spectrum. We investigate several explanations for this feature, including thermal emission due to -process nucleosynthesis, nebular Fe lines, and He emission, but cannot definitively identify its origin with present information.
We fit SN 2025kg with a one-zone radioactive decay model and a model that accounts for mixing of 56Ni into the outer SN ejecta. We derive a range in the 56Ni mass of .
Together, observations presented in our companion paper (Eyles-Ferris et al., 2025) and this work favor a ZAMS progenitor mass for EP 250108a/SN 2025kg of or if the compact remnant is a neutron star or black hole, respectively. This progenitor was surrounded by a He shell ejected during a common envelope episode, and either H-rich clumps due to an prior, asymmetric common envelope ejection or a giant/supergiant tertiary stellar companion.
Our analysis supports a causal link between GRB-SNe and FXT SNe, in which GRBs are produced by successful jets and FXT SNe are produced by trapped or weak jets. Thus far, the rate of SN detections following EP FXTs indicates that trapped or weak jets are substantially more common than successful (GRB) jets.
EP 250108a is only the second EP transient with an observed SN counterpart to date. Our detailed observational study of SN 2025kg has revealed both similarities with previous FXT and GRB SNe, and unexpected signals, including H at 42.5 days and the broad 4-4.5m feature, that cannot be fully accounted for with current models. Looking forward, the opportunities for detailed observational studies to shed light on the stellar progenitors of FXTs and their catastrophic explosions is only increasing with the continued operations of EP and coordinated rapid follow-up efforts.
Acknowledgements
We dedicate this work to the memory of Alicia M. Soderberg, a pioneer in the study of XRFs and FXTs.
We are deeply grateful to Tom Marsh for developing the molly software, one of his many contributions to advancing the field of compact objects. We thank Jennifer Andrews, the T-80 South technical team, and others for their support during observations.
J.C.R. acknowledges support from the Northwestern Presidential Fellowship. P.G.J. J.N.D.D., J.S.S., J.Q.V., and A.P.C.H. are supported by the European Union (ERC, Starstruck, 101095973). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. N.S. acknowledges support from the Knut and Alice Wallenberg Foundation through the “Gravity Meets Light” project and by and by the research environment grant “Gravitational Radiation and Electromagnetic Astrophysical Transients” (GREAT) funded by the Swedish Research Council (VR) under Dnr 2016-06012. B.P.G acknowledges support from STFC grant No. ST/Y002253/1 and The Leverhulme Trust grant No. RPG-2024-117 C.D.K. gratefully acknowledges support from the NSF through AST-2432037, the HST Guest Observer Program through HST-SNAP-17070 and HST-GO-17706, and from JWST Archival Research through JWST-AR-6241 and JWST-AR-5441. W.F. gratefully acknowledges support by the David and Lucile Packard Foundation, the Alfred P. Sloan Foundation, and the Research Corporation for Science Advancement through Cottrell Scholar Award 28284. P.O.’B. acknowledges support from the UKRI grant: ST/W000857/1. The work by C.L.F. was supported by the US Department of Energy through the Los Alamos National Laboratory. Los Alamos National Laboratory is operated by Triad National Security, LLC, for the National Nuclear Security Administration of U.S. Department of Energy (Contract No. 89233218CNA000001). W.J-G. is supported by NASA through the NASA Hubble Fellowship grant HSTHF2-51558.001-A awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., for NASA, under contract NAS5-26555. L.M.R acknowledges support from NSF grants AST-1911140, AST-1910471, and AST-2206490. S.Y. acknowledges the funding from the National Natural Science Foundation of China under grant No.12303046 and from the Startup Research Fund of Henan Academy of Sciences No.241841217. A.A.C. acknowledges support through the European Space Agency (ESA) research fellowship programme. C.R.B. acknowledges the financial support from CNPq (316072/2021-4) and from FAPERJ (grants 201.456/2022 and 210.330/2022) and the FINEP contract 01.22.0505.00 (ref. 1891/22). D.M.S. and M.A.P.T. acknowledge support by the Spanish Ministry of Science via the Plan de Generacion de conocimiento PID2020-120323GB-I00. DMS also acknowledges support via a Ramon y Cajal Fellowship RYC2023-044941. M.N. is supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 948381).. R.G. was sponsored by the National Aeronautics and Space Administration (NASA) through a contract with ORAU. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the National Aeronautics and Space Administration (NASA) or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein. S.J.S., J.G., S.S. acknowledge funding from STFC Grant ST/Y001605/1, a Royal Society Research Professorship and the Hintze Charitable Foundation. C.J.N. acknowledges support from the Science and Technology Facilities Council (grant No. ST/Y000544/1) and from the Leverhulme Trust (grant No. RPG-2021-380). A.A. acknowledges the Yushan Young Fellow Program by the Ministry of Education, Taiwan for the financial support (MOE-111-YSFMS-0008-001-P1). T.-W.C. acknowledges the Yushan Fellow Program by the Ministry of Education, Taiwan for the financial support (MOE-111-YSFMS-0008-001-P1). C.L. is supported by DoE award #DE-SC0025599. A.R.E. is supported by the European Space Agency (ESA) Research Fellowship.
Observations reported here were obtained at the MMT Observatory, a joint facility of the University of Arizona and the Smithsonian Institution. MMT Observatory access was supported by Northwestern University and the Center for Interdisciplinary Exploration and Research in Astrophysics (CIERA).
Based on observations obtained at the international Gemini Observatory (Program IDs GN-2024B-Q-105, GN-2024B-Q-107, GS-2024B-Q-105, GS-2025A-Q-107), a program of NOIRLab, which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation on behalf of the Gemini Observatory partnership: the National Science Foundation (United States), National Research Council (Canada), Agencia Nacional de Investigación y Desarrollo (Chile), Ministerio de Ciencia, Tecnología e Innovación (Argentina), Ministério da Ciência, Tecnologia, Inovações e Comunicações (Brazil), and Korea Astronomy and Space Science Institute (Republic of Korea). Data was processed using the Gemini DRAGONS (Data Reduction for Astronomy from Gemini Observatory North and South) package.
This work is based in part on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with program #6133.
Data for this paper has in part been obtained under the International Time Programme of the CCI (International Scientific Committee of the Observatorios de Canarias of the IAC) under programm ID ITP24 PI Jonker with the NOT and GTC operated on the island of La Palma by the Roque de los Muchachos. Observations have been made in part with the ALFOSC instrument, which is provided by the Instituto de Astrofisica de Andalucia (IAA) under a joint agreement with the University of Copenhagen and the Nordic Optical Telescope, owned in collaboration by the University of Turku and Aarhus University, and operated jointly by Aarhus University, the University of Turku and the University of Oslo, representing Denmark, Finland and Norway, the University of Iceland and Stockholm University at the Observatorio del Roque de los Muchachos, La Palma, Spain, of the Instituto de Astrofisica de Canarias.
This publication has made use of data collected at Lulin Observatory, partly supported by MoST grant 109-2112-M-008-001 and TAOVA with NSTC grant 113-2740-M-008-005.
Based on observations obtained with MegaPrime/MegaCam, a joint project of CFHT and CEA/DAPNIA, at the Canada-France-Hawaii Telescope (CFHT) which is operated by the National Research Council (NRC) of Canada, the Institut National des Science de l’Univers of the Centre National de la Recherche Scientifique (CNRS) of France, and the University of Hawai’i. The observations at the Canada-France-Hawaii Telescope were performed with care and respect from the summit of Maunakea which is a significant cultural and historic site.
Based on observations with the BlackGEM telescope array. The BlackGEM telescope array is built and run by a consortium consisting of Radboud University, the Netherlands Research School for Astronomy (NOVA), and KU Leuven with additional support from Armagh Observatory and Planetarium, Durham University, Hamburg Observatory, Hebrew University, Las Cumbres Observatory, Tel Aviv University, Texas Tech University, Technical University of Denmark, University of California Davis, the University of Barcelona, the University of Manchester, University of Potsdam, the University of Valparaiso, the University of Warwick, and Weizmann Institute of science. BlackGEM is hosted and supported by ESO at La Silla.
Pan-STARRS is primarily funded to search for near-Earth asteroids through NASA grants NNX08AR22G and NNX14AM74G. The Pan-STARRS science products for transient follow-up are made possible through the contributions of the University of Hawaii Institute for Astronomy and Queen’s University Belfast.
This work has made use of data from the Asteroid Terrestrial-impact Last Alert System (ATLAS) project. The Asteroid Terrestrial-impact Last Alert System (ATLAS) project is primarily funded to search for near earth asteroids through NASA grants NN12AR55G, 80NSSC18K0284, and 80NSSC18K1575; byproducts of the NEO search include images and catalogs from the survey area. This work was partially funded by Kepler/K2 grant J1944/80NSSC19K0112 and HST GO-15889, and STFC grants ST/T000198/1 and ST/S006109/1. The ATLAS science products have been made possible through the contributions of the University of Hawaii Institute for Astronomy, the Queen’s University Belfast, the Space Telescope Science Institute, the South African Astronomical Observatory, and The Millennium Institute of Astrophysics (MAS), Chile.
The Legacy Surveys consist of three individual and complementary projects: the Dark Energy Camera Legacy Survey (DECaLS; Proposal ID #2014B-0404; PIs: David Schlegel and Arjun Dey), the Beijing-Arizona Sky Survey (BASS; NOAO Prop. ID #2015A-0801; PIs: Zhou Xu and Xiaohui Fan), and the Mayall -band Legacy Survey (MzLS; Prop. ID #2016A-0453; PI: Arjun Dey). DECaLS, BASS and MzLS together include data obtained, respectively, at the Blanco telescope, Cerro Tololo Inter-American Observatory, NSF’s NOIRLab; the Bok telescope, Steward Observatory, University of Arizona; and the Mayall telescope, Kitt Peak National Observatory, NOIRLab. Pipeline processing and analyses of the data were supported by NOIRLab and the Lawrence Berkeley National Laboratory (LBNL). The Legacy Surveys project is honored to be permitted to conduct astronomical research on Iolkam Duág (Kitt Peak), a mountain with particular significance to the Tohono Oódham Nation.